ABC is a given triangle. Forces P,Q,R acting along the lines OA,CB,OC are in equilibrium. Show that P:Q:R=a^2(b^2+c^2-a^2):b^2(c^2+a^2-b^2):c^2(a^2+b^2-c^2) if O is the circumcentre of the triangle

Grade:12th pass

2 Answers

Srutarshi Tripathi
3 years ago
By Lami‟s theorem, P/sin(BOC) = Q/sin(COA) = R/sin(AOB)OR, P/sin(2A) = Q/sin(2B) = R/sin(2C)Or, P/2sinA.cosA = Q/2sinB.cosB = R/2sinC.cosC ………………………….(1)WE HAVE, cos A = (b^2 + c^2 – a^2)/2bcsinA = 2.(area)/bcso, 2sinAcosA = 2(2.(area)/bc).( (b^2 + c^2 – a^2)/2bc) =2.(area)( b^2 + c^2 – a^2)/b^2 .c^2Similarly we find 2sinB.cosB and 2sinC.cosC and putting these values in (1) and dividing it by a^2.b^2.c^2, we get the required result
Siva
3 years ago
A body is at rest on a rough inclined plane of inclination 𝛼 to the horizon, being acted on
by a force making an angle 𝜃 with the plane. Find the limit between which the force
must lie and also find the magnitude and direction of the least force required to drag the
body up the inclined plane